OpenStudy (fanduekisses):

How to verify the identity tan(x+pi/2)=-cotx ?

OpenStudy (fanduekisses):

DoI use the cofunction identiy first?

OpenStudy (whpalmer4):

Well, one way would be to expand this in terms of the basic functions: \[\tan x = \frac{\sin x}{\cos x}\]\[\cot x = \frac{\cos x}{\sin x}\]and work through it using the identities for the sin and cos of a sum